# IberFire - a detailed creation of a spatio-temporal dataset for wildfire risk assessment in Spain

### v1

@@ Abstract
Wildfires pose a critical environmental issue to ecosystems, economies, and public safety, particularly in Mediterranean regions such as Spain. Accurate predictive models rely on high-resolution spatio-temporal data to capture the complex interplay of environmental and anthropogenic factors. To address the lack of localised and fine-grained datasets in Spain, this work introduces IberFire, a spatio-temporal datacube at 1 km × 1 km × 1-day resolution covering mainland Spain and the Balearic Islands from December 2007 to December 2024. IberFire integrates 260 features across eight main categories: auxiliary features, fire history, geography, topography, meteorology, vegetation indices, human activity, and land cover. All features are derived from open-access sources, ensuring transparency and real-time applicability. The data processing pipeline was implemented entirely using open-source tools, and the codebase has been made publicly available. This work not only enhances spatio-temporal granularity and feature diversity compared to existing European datacubes but also provides a reproducible methodology for constructing similar datasets. IberFire supports advanced wildfire risk modelling through Machine Learning (ML) and Deep Learning (DL) techniques, enables climate pattern analysis and informs strategic planning in fire prevention and land management. The dataset is publicly available on Zenodo to promote open research and collaboration.
@@ 1 Background & Summary
Forest fires constitute a critical environmental issue with severe ecological, social, and economic implications. Wildfires not only destroy vast forest areas, cause the loss of natural habitats, and release large amounts of carbon dioxide; they can also cause substantial economic damage through the destruction of infrastructure, housing, and productive land.
Spain is one of the most affected countries within the European Union [1, 2]. Nearly 40% of the total burned area in the entire Mediterranean region of Europe between 1980 and 2008 was in Spain [3]. Furthermore, data from the European Forest Fire Information System (EFFIS) [4] indicate that over 7,000 fires have occurred in Spain since 2008, as can be seen in the left image of Figure 1.
@@
Wildfires are progressively increasing in scale, with a growing number of autonomous communities experiencing wildfires spanning areas of 5,000 ha, 10,000 ha, and even 20,000 ha [5]. The year 2022 marked the most severe wildfire season, with two consecutive forest fires jointly burning over 60,000 ha, an area roughly equivalent to that of Madrid.
@@
In this context, the development of precise predictive models can help with prevention and fire service management by providing early warnings and identifying high-risk areas. Physical models, such as the Canadian Fire Weather Index [6], leverage meteorological data and physical equations to make predictions; however, recent studies have shown that data-driven models often achieve superior performance in terms of predictive accuracy [7, 8].
@@
Building Machine Learning (ML) and Deep Learning (DL) models for fire risk assessment requires high-resolution spatio-temporal features. Many data sources are available for this purpose: the Corine Land Cover (CLC) [9] dataset provides a classification of the land usage, ERA5-Land [10] offers hourly meteorological curated data and vegetation indices can be retrieved from the Copernicus Land Monitoring Service (CLMS) [11]. When complemented with additional variables, these data enable the development of robust fire-risk prediction models. However, these sources differ significantly in spatial and temporal resolution, format, and update frequency, making direct integration a non-trivial task.
@@
Datacubes are multidimensional data structures designed to standardise spatial and spatio-temporal features with varying original resolutions, providing a consistent and accessible means of analysis. They facilitate the modelling of complex spatio-temporal phenomena without the need for separate pipelines for each data source. In the context of wildfire risk prediction, datacubes are particularly crucial, as they enable the integration of heterogeneous environmental variables that influence fire behaviour, like the CLC dataset and ERA5-Land, alongside historical fire records.
To the best of current knowledge, only two datacubes that include Spain within their area of interest are available for this purpose, although neither is specifically focused on the Spanish territory. On the one hand, SeasFire Cube [12] offers 59 variables from 2001 to 2021 with 0.25°$\degree$ spatial resolution (approximately 27 km at the equator) and 8-day temporal resolution. While this dataset may be used for fire-risk predictions, it is suggested that its resolution is not sufficiently granular for practical applicability at the scale of Spain. On the other hand, Mesogeos [13] offers a 1 km spatial resolution covering the Mediterranean area from 2006 to 2022, with 27 spatio-temporal features. In this case, it is believed that incorporating a broader range of Spain-specific features could enhance forest fire risk predictive models, potentially improving the accuracy of the predictions.
@@
The IberFire [14] datacube was constructed to address this gap, a 1km $\times$ 1km $\times$ 1-day high-resolution datacube covering Spain from December 2007 to December 2024 with 260 different features related to forest fire risk. All features were selected based on their potential to be automatically retrieved from external sources, allowing for real-time model deployment.
@@
The features of IberFire can be divided into 8 main categories: auxiliary features that help locate the cells, fire history obtained from EFFIS, geographical location information features, land usage from Copernicus Corine Land Cover [9], topography variables obtained from the European Digital Elevation Model [15], human activity related features retrieved from WorldPop[16] and OpenStreetMap [17], meteorological variables obtained from ERA5-Land [10], and vegetation indices downloaded from Copernicus Land Monitoring Service [11].
@@
This paper presents two main objectives. The first is the introduction and public release of the IberFire datacube, which improves upon existing datasets in terms of resolution and feature diversity for Spain. The IberFire datacube offers high-resolution modelling capabilities to gain insights not only into Spain’s fire risk behaviour but also into climate-change pattern time-series modelling. The second objective is the provision of a reproducible, systematic methodology for constructing similar datacubes, an approach that can be extended to model other spatio-temporal environmental phenomena. This includes a detailed explanation of the generation of IberFire, along with the concepts needed to manipulate geospatial data.
@@ 1.1 Concepts and tools about Geographic Information Systems (GIS)
Geographic features are not usually stored in commonly used formats such as CSV (Comma Separated Values); instead, specific formats that contain integrated geographical coordinates, such as raster data, are used. Among these, datacubes organise spatial and temporal information into structured, multi-dimensional arrays. This structure enables efficient storage, retrieval, and analysis of geospatial variables over both space and time, since any cell of the grid can be accessed and every feature value of that cell can be retrieved. The construction of such a datacube requires an understanding of Geographic Information Systems (GIS). This section provides an introduction to GIS, emphasising the main data formats and processing techniques involved in manipulating spatial data.
@@
Vector data represents geographic features using points, lines, and polygons, which accurately capture geometric locations and boundaries [18]. Each instance of a vector dataset corresponds to a geographic shape with some feature values assigned. This representation is particularly suited for discrete features such as roads or specific fire-affected areas, as exemplified in the left image of Figure 2.
Figure 2: Left: Example of vectorial data (blue), some burned areas retrieved from EFFIS. Right: Example of raster data, elevation values on the same region as the left plot at a 1km×$\times$1km resolution.
On the other hand, raster data represents geographic space as a regular grid of cells [19], where each pixel is assigned a value corresponding to a property of the geographic area, as can be seen in the right image of Figure 2. This data format is commonly used to represent continuous geographic phenomena such as climate data.
@@
Effectively processing geospatial data often requires the use of specialised tools. QGIS [20] is an open-source software employed for working with vector and raster data. This software offers an extensive set of functions for visualising, manipulating, automatically reprojecting, and resampling spatial datasets. Other tools for working with GIS data include the rasterio and xarray Python libraries, which provide modules for resampling raster data and creating datacubes, respectively. In this work, both QGIS and Python-based tools were employed, ensuring the reproducibility of the entire process through open-source software solutions.
@@
As some features do not vary over time, in order not to store repeated values, they can be saved as spatial-only features; hence, only one array for the entire period that the datacube covers. For instance, elevation values could be stored as spatial-only features. Figure 3 provides a visual representation of the datacube concept.
@@ 2 Methods
This chapter provides a detailed analysis of the construction process of the IberFire datacube. Section 2.1 describes the generation of the spatio-temporal grid. After that, Section 2.2 analyses the auxiliary features introduced to the datacube. Then, Section 2.3 presents a thorough explanation of the integration of the primary output feature, is_fire. Subsequently, Sections 2.4 to 2.9 describe the incorporation of the explanatory variables into the dataset. Finally, Section 2.10 provides a summary of all external sources utilized in the construction of the datacube.
@@ 2.1 Grid generation
Given that datacubes are many equal-shaped raster arrays arranged along the time dimension, the grid of a datacube is required to be rectangular. A key consideration in the creation of the grid was ensuring that each cell represented exactly 1 km2 of area. To achieve this, the grid was created on the EPSG:3035 CRS, whose units are meters. This approach ensures that all cells have the same importance and no cell is underrepresented. Within QGIS, the region of interest shown in the right image of Figure 1 was selected, and an empty base raster file with a 1km $\times$ 1km spatial resolution was generated on that region. This empty raster file was then saved as a reference for the subsequent generation of the datacube and the creation of all the layers it contains.
The datacube was then constructed in Python using the xarray package. To build a datacube, it is necessary to define coordinates in each dimension, hence, for $x$, $y$, and $t$ dimensions. Each coordinate consists of a vector of values that are used to locate the individual grid cells. These coordinates are similar to indices, but they can be any ordered set of values, including temporal sequences such as dates.
The base raster file’s horizontal and vertical coordinate values served as the coordinates of the $x$ and $y$ dimensions of the datacube respectively. Then, for the temporal dimension, the coordinates consisted of daily timestamps from 01/12/2007 to 31/12/2024. The final datacube comprised 1188 values for the $x$ dimension, 920 values for the $y$ dimension, and 6241 values for the time axis. As a result, the datacube contains a total of $1188\cdot 920\cdot 6241\approx 6.8\cdot 10^{9}$ different cells, each storing different features.
@@ 2.2 Auxiliary features
After constructing the empty datacube, three auxiliary features were added to the dataset: x_index, y_index, and is_spain. These features, as analysed in Chapter 3, are not intended to serve as explanatory variables, but instead facilitate the manipulation and processing of the datacube.
@@
The third auxiliary feature, is_spain, identifies the cells corresponding to Spanish territory, which are the only ones for which predictions should be generated. This layer was derived from a vectorial dataset containing the boundaries of Spain obtained from simplemaps [21] and processed with the QGIS software.
@@ 2.3 Fire history: EFFIS
The historical fire data for Spain were obtained using the EFFIS data request format [4]. The raw data were retrieved in vectorial format and contained geometries representing the burned area of historical fire events, along with the corresponding start and end dates of each fire event. To integrate this information into the datacube, the intersection of the spatial grid cells with the fire geometries in QGIS was calculated, as illustrated in Figure 4. Subsequently, a binary layer in the datacube, named is_fire, was created. A value of 1 was assigned to cells that intersected with fire geometries and fell within the corresponding temporal interval defined by the fire’s start and end dates. Following the definition of fire danger from [7], the previous process was executed on fires (geometries retrieved from EFFIS) with a burned area greater than 5 ha. This is because fire danger can be viewed as the combined risk of a fire igniting and the risk of that fire growing large (>5 ha). Introducing small wildfires as fire instances (is_fire = 1) could lead to inconsistencies, as these small fires did not grow larger for certain reasons (for example, high humidity). Therefore, the fire risk for these small fires was low, even though a fire was present.
@@
After that, the is_near_fire layer was added, which is a binary feature indicating whether a cell is within a 25 $\times$ 25 cell area centred on each is_fire = 1 instance, as well as the 10 days preceding the event. This results in a 25 $\times$ 25 $\times$ 10-sized box prior to each fire cell. This feature is useful for analysing the spatial and temporal proximity of fire occurrences, enabling the investigation of both the geographic and temporal closeness of future fires in relation to past events.
@@ 2.4 Geographical location information
The first two features consist of the values of the coordinates of each cell in the EPSG:3035 coordinate reference system. The next two features, is_sea and is_waterbody, are binary indicators that denote whether a given cell is located over open sea or inland water, respectively. These features were calculated using QGIS with data from the European Digital Elevation Model.
Lastly, the AutonomousCommunities feature represents the level 2 NUTS (Nomenclature of Territorial Units for Statistics) [22] division of Spain, with values listed in Table 1. However, since the region of interest does not include Ceuta, Melilla, and the Canary Islands, the corresponding values for these regions do not appear in the dataset. The feature was generated in QGIS by converting a vector dataset containing the shapes of the autonomous communities of Spain [23] into a raster layer, assigning to each cell the corresponding value from Table 1 based on its intersection with the appropriate autonomous community.

### v2

@@ Abstract
Wildfires pose a threat to ecosystems, economies and public safety, particularly in Mediterranean regions such as Spain. Accurate predictive models require high-resolution spatio-temporal data to capture complex dynamics of environmental and human factors. To address the scarcity of fine-grained wildfire datasets in Spain, we introduce IberFire: a spatio-temporal dataset with 1 km × 1 km × 1-day resolution, covering mainland Spain and the Balearic Islands from December 2007 to December 2024. IberFire integrates 120 features across eight categories: auxiliary data, fire history, geography, topography, meteorology, vegetation indices, human activity and land cover. All features and processing rely on open-access data and tools, with a publicly available codebase ensuring transparency and applicability. IberFire offers enhanced spatial granularity and feature diversity compared to existing European datasets, and provides a reproducible framework. It supports advanced wildfire risk modelling via Machine Learning and Deep Learning, facilitates climate trend analysis, and informs fire prevention and land management strategies. The dataset is freely available on Zenodo to promote open research and collaboration.
@@ 1 Background & Summary
Forest fires constitute a critical environmental issue with severe ecological, social, and economic implications. Wildfires not only destroy vast forest areas, cause the loss of natural habitats, and release large amounts of carbon dioxide, but also cause substantial economic damage through the destruction of infrastructure, housing, and productive land.
Spain is one of the countries most affected within the European Union [rodrigues_insight_2014, oliveira_modeling_2012]. Nearly 40% of the total burned area in the entire Mediterranean region of Europe between 1980 and 2008 was in Spain [moreno2012modelo]. Furthermore, data from the European Forest Fire Information System (EFFIS) [effis] indicate that over 7,000 fires have occurred in Spain since 2008, as shown in the left image of Figure 1.
@@
Wildfires are increasing in scale, with a growing number of autonomous communities experiencing wildfires spanning areas of 5,000 ha, 10,000 ha, and even 20,000 ha [ubeda_grandes_2023]. The year 2022 marked the most severe wildfire season, with two consecutive forest fires jointly burning over 60,000 ha in the same region, an area roughly equivalent to that of Madrid.
@@
In this context, the development of precise predictive models can support fire prevention and fire service management by providing early warnings and identifying high-risk areas. Physical models, such as the Canadian Fire Weather Index [FWI], leverage meteorological data and physical equations to make predictions; however, recent studies have shown that data-driven models often achieve superior performance in terms of predictive accuracy [prapas_deep_2021, kondylatos_wildfire_2022].
@@
Building Machine Learning (ML) and Deep Learning (DL) models for fire risk assessment requires high-resolution spatio-temporal features. Many data sources are available for this purpose: the Corine Land Cover (CLC) [CLC] dataset provides a classification of land usage, ERA5-Land [ERA5_validation] offers hourly meteorological curated data and vegetation indices can be retrieved from the Copernicus Land Monitoring Service (CLMS) [CopernicusLandMonitoringService]. When complemented with additional variables, these data enable the development of robust fire-risk prediction models. However, these sources differ significantly in spatial and temporal resolution, format, and update frequency, making direct integration a non-trivial task.
@@
Datacubes are multidimensional data structures designed to standardise spatial and spatio-temporal features with varying original resolutions, providing a consistent and accessible means of analysis. They facilitate the modelling of complex spatio-temporal phenomena eliminating the need for independent processing pipelines for each data source. In the context of wildfire risk prediction, datacubes are particularly crucial, as they enable the integration of historical fire records alongside heterogeneous environmental variables that influence fire behaviour, like the CLC dataset and ERA5-Land.
[…]
@@
The IberFire [erzibengoa_2025_15225886] datacube was constructed to address this gap. It is a 1km $\times$ 1km $\times$ 1-day high-resolution datacube covering Spain from December 2007 to December 2024. It includes 120 features identified in the literature as relevant to forest fire risk. All features were selected based on their potential to be automatically retrieved from external sources, allowing for real-time model deployment.
@@
The features of IberFire can be divided into 8 main categories: auxiliary features that assist in locating the cells, fire history obtained from EFFIS, geographical location information features, land usage from Copernicus Corine Land Cover [CLC], topography variables obtained from the European Digital Elevation Model [EU_DEM], human activity related features retrieved from WorldPop[worldpop] and OpenStreetMap [openstreetmap], meteorological variables obtained from ERA5-Land [ERA5_validation], and vegetation indices downloaded from Copernicus Land Monitoring Service [CopernicusLandMonitoringService].
@@
This paper presents two main objectives. The first is the introduction and public release of the IberFire datacube, which improves upon existing datasets in terms of resolution and feature diversity for Spain. The IberFire datacube offers high-resolution modelling capabilities to gain insights not only into Spain’s fire risk behaviour but also into time-series modelling of climate change patterns. The second objective is the provision of a reproducible, systematic methodology for constructing similar datacubes, an approach that can be extended to model other spatio-temporal environmental phenomena. The presented methodology includes a detailed explanation of the generation of IberFire, along with the concepts needed to manipulate geospatial data.
@@ 1.1 Concepts and tools about Geographic Information Systems (GIS)
Geographic features are not usually stored in commonly used formats such as CSV (Comma Separated Values); instead, specific formats that contain integrated geographical coordinates, are used; raster data is an example of this. Among these formats, datacubes organise spatial and temporal information into structured, multi-dimensional arrays. This structure enables efficient storage, retrieval, and analysis of geospatial variables over both space and time, since any cell of the grid can be accessed and every feature value of that cell can be retrieved. The construction of such a datacube requires an understanding of Geographic Information Systems (GIS). This subsection provides an introduction to GIS, emphasising the main data formats and processing techniques involved in manipulating spatial data.
@@
Vector data represents geographic features using points, lines, and polygons, which accurately capture geometric locations and boundaries [QGIS_VectorAttributeData]. Each instance of a vector dataset corresponds to a geographic shape with some feature values assigned. This representation is particularly suited for discrete features such as roads or specific fire-affected areas, as exemplified in the left image of Figure 2.
Figure 2: Left: Example of vectorial data (blue), some burned areas retrieved from EFFIS. Right: Example of raster data, elevation values on the same region as the left plot at a $\times1km$ resolution.
On the other hand, raster data represents geographic space as a regular grid of cells [QGIS_raster_data], where each pixel is assigned a value corresponding to a property of the geographic area, as can be seen in the right image of Figure 2. This data format is commonly used to represent continuous geographic phenomena such as climate data.
@@
Effectively processing geospatial data often requires the use of specialised tools. QGIS [QGIS] is an open-source software employed for working with vector and raster data. This software offers an extensive set of functions for visualising, manipulating, automatically reprojecting, and resampling spatial datasets. Other tools for working with GIS data include the rasterio and xarray Python libraries, which provide modules for resampling raster data and creating datacubes, respectively. In this work, both QGIS and Python-based tools were employed, ensuring the reproducibility of the entire process through open-source software solutions.
@@
As some features do not vary over time, in order not to store repeated values, they can be saved as spatial-only features; hence, only one array is saved for the entire period that the datacube covers. For instance, elevation values could be stored as a spatial-only feature, whereas temperature values should be stored as spatio-temporal. Figure 3 provides a visual representation of the datacube concept.
@@
This storage approach is particularly useful for features with low temporal update frequency. For example, features that update annually—although technically spatio-temporal—can be stored as spatial-only layers. This avoids storing 365 identical rasters per year and instead requires just a single raster array, significantly reducing redundancy and storage requirements.
@@ 2 Methods
This section provides a detailed analysis of the construction process of the IberFire datacube. Subsection 2.1 describes the generation of the spatio-temporal grid. After that, Subsection 2.2 analyses the auxiliary features introduced to the datacube. Then, Subsection 2.3 presents a thorough explanation of the integration of the primary output feature, is_fire, alongside the integration of the baseline model, the Fire Weather Index (FWI). Subsequently, Subsections 2.4 to 2.9 describe the incorporation of the explanatory variables into the dataset. Finally, Subsection 2.10 provides a summary of all external sources utilised in the construction of the datacube.
@@ 2.1 Grid generation
Given that datacubes are many equal-shaped raster arrays arranged along the time dimension, the grid of a datacube is required to be rectangular. A key consideration in the creation of the grid was ensuring that each cell represented exactly 1 km2 of area, which ensures that all cells have the same importance and no cell is underrepresented. To achieve this, the grid was created on the EPSG:3035 CRS, whose units are metres. Within QGIS, the region of interest shown in the right image of Figure 1 was selected, and an empty base raster file with a 1km $\times$ 1km spatial resolution was generated on that region. This empty raster file was then saved as a reference for the subsequent generation of the datacube and the creation of all the layers it contains.
The datacube was then constructed in Python using the xarray package. To build a datacube, it is necessary to define coordinate values in each dimension, hence, for $x$, $y$, and $t$ dimensions. Each coordinate consists of a vector of values that are used to locate the individual grid cells. These coordinates are similar to indices, but they can be any ordered set of values, including temporal sequences such as dates.
The base raster file’s horizontal and vertical coordinate values served as the coordinates of the $x$ and $y$ dimensions of the datacube respectively. Then, for the temporal dimension, the coordinates consisted of daily timestamps from 01/12/2007 to 31/12/2024. The final datacube comprised 1188 values for the $x$ dimension, 920 values for the $y$ dimension, and 6241 values for the time axis. As a result, the datacube contains a total of $1188\cdot 920\cdot 6241\approx 6.8\cdot 10^{9}$ different cells, each storing different features. However, not all of these cells should be used for modelling, since not all the cells fall inside Spain, as mentioned in Section 5. The amount of usable cells is $498530\cdot 6241\approx 3.1\cdot 10^{9}$.
@@ 2.2 Auxiliary features
After constructing the empty datacube, three auxiliary features were added to the dataset: x_index, y_index, and is_spain. These features, as explained in more detail in Section 5, are not intended to serve as explanatory variables, but rather to facilitate the manipulation and processing of the datacube.
@@
The third auxiliary feature, is_spain, identifies the cells corresponding to Spanish territory, which are the only ones for which predictions should be generated. This layer was derived from a vectorial dataset containing the boundaries of Spain obtained from simplemaps [simplemaps-spain] and processed with the QGIS software.
@@ 2.3 Fire history: EFFIS
The historical fire data for Spain were obtained using the EFFIS data request format [effis]. The raw data were retrieved in vectorial format and contained geometries representing the burned area of historical fire events, along with the corresponding start and end dates of each fire event. To integrate this information into the datacube, the intersection of the spatial grid cells with the fire geometries in QGIS was calculated, as illustrated in Figure 4. Subsequently, a binary layer in the datacube, named is_fire, was created. A value of 1 was assigned to cells that intersected with fire geometries and fell within the corresponding temporal interval defined by the fire’s start and end dates. Following the definition of fire danger from [prapas_deep_2021], the previous process was executed on fires (geometries retrieved from EFFIS) with a burned area greater than 5 ha. This is because fire danger can be viewed as the combined risk of a fire igniting and the risk of that fire growing large (>5 ha). Introducing small wildfires as fire instances (is_fire = 1) could lead to inconsistencies, as these small fires did not grow larger for certain reasons (for example, high humidity). Therefore, the fire risk for these small fires was low, even though a fire was present.
@@
After that, the is_near_fire layer was added, which is a binary feature indicating whether a cell is within a 25 $\times$ 25 cell area centred on each is_fire = 1 instance, as well as the 10 days preceding the event. This results in a 25 $\times$ 25 $\times$ 10-sized box prior to each fire cell. The is_near_fire feature is particularly useful for identifying true non-fire events that are neither spatially nor temporally close to fire events, thereby avoiding the inclusion of near-fire instances that may closely resemble actual fire conditions due to their proximity.
@@
Baseline model: Fire Weather Index
The Fire Weather Index (FWI) [FWI] was introduced to IberFire as a baseline model. It leverages temperature, wind, relative humidity and precipitation data, along with physical equations to predict a continuous value that represents fire risk. The FWI takes values in the range $0,+\inf)$, although the most common values lie between 0 and 50. Depending on the value that it has, fire risk levels are assigned according to Table 1 [FWI_explanations].
Very low Low Moderate Hight Very high Extreme
<< 5.2 5.2 - 11.2 11.2 - 21.3 21.3 -38.0 38.0 - 50 $\geq$ 50
Table 1: FWI risk levels.
To introduce FWI to IberFire, data from the Copernicus Emergency Management Service (CEMS) [FWI_data] was downloaded, which provides daily values at a spatial resolution of 0.25º × 0.25º latitude-longitude, equivalent to approximately 27.5 km. Since the original dataset resolution is lower than desired, the data was interpolated to a 1 km × 1 km resolution using nearest-neighbor interpolation. This feature was also added as a spatio-temporal feature.
@@ 2.4 Geographical location information
The first two features consist of the values of the coordinates of each cell in the EPSG:3035 coordinate reference system. The next two features, is_sea and is_waterbody, are binary indicators that denote whether a given cell is located over open sea or inland water, respectively. These features were calculated using QGIS with data from the European Digital Elevation Model [EU_DEM].
Lastly, the AutonomousCommunities feature represents the level 2 NUTS (Nomenclature of Territorial Units for Statistics) [nuts_ec_2016] division of Spain, with values listed in Table 2. However, since the region of interest does not include Ceuta, Melilla, and the Canary Islands, the corresponding values for these regions do not appear in the dataset. The feature was generated in QGIS by converting a vector dataset containing the shapes of the autonomous communities of Spain [area2018limites] into a raster layer, assigning to each cell the corresponding value from Table 2 based on its intersection with the appropriate autonomous community.